OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form

OpenStudy (opcode):

Use this formula.

OpenStudy (anonymous):

\[\frac{ -3-6 }{ 0-3 }\]

OpenStudy (anonymous):

slope is 3

OpenStudy (opcode):

Here's another example: I used: "(0,0), (3,-6)" \(\huge m =\frac {y^2 - y^1}{x^2 - x^1} = \frac {-6 - 0}{3 - 0} = \frac {-6}{3} = -2\)

OpenStudy (anonymous):

opcode....

OpenStudy (anonymous):

YOUR WRONG!!!

OpenStudy (anonymous):

WTF??

OpenStudy (anonymous):

thats not the slope

OpenStudy (anonymous):

using that same formula

OpenStudy (anonymous):

you just placed the digits in the wrong place

OpenStudy (anonymous):

\[m=\frac{ -3-6 }{ 3-0 }=\frac{ -9 }{ 3 }=-3\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

sorry. its 0-3 = -3

OpenStudy (anonymous):

final answer= 3

OpenStudy (opcode):

Sorry about that. Haven't slept in a long time :S